OpenStudy (anonymous):

A farmer has a rectangular field which measures 100 feet for the length and w feet for the width. If the perimeter is the sum of all four sides of the field, what is an expression for the perimeter of the field? Simplify

OpenStudy (anonymous):

P=2*100+2w=200+2w

OpenStudy (anonymous):

100w or 200 + 2w

OpenStudy (anonymous):

200+w^2

OpenStudy (anonymous):

perimeter = 200 +2w that is 2*length + 2 * w

OpenStudy (anonymous):

w + w = 2w not w^2

OpenStudy (anonymous):